OpenStudy (anonymous):

Write an equation of an ellipse in standard form with the center at the origin and with the given characteristics. vertex at (3, 0) and co-vertex at (0, 5)

OpenStudy (tkhunny):

Well, do it! \(\dfrac{(x - h)^{2}}{a^{2}} + \dfrac{(y-k)^{2}}{b^{2}} = 1\) Fill in the blanks!

OpenStudy (anonymous):

i have no idea how

OpenStudy (anonymous):

its mutiply choice

OpenStudy (tkhunny):

It makes no difference what kind of problem statement is presented. If you truly have "no idea", then you cannot solve this problem. You need to read your book, attend class, and come back with the right information. Sorry.